# Gluon Distributions in the Pion

### v1

@@ I Introduction
Global analyses of pion PDFs, primarily using Drell-Yan, $J/\psi$ production, and direct photon data Barry et al. (2021); Novikov et al. (2020); Pasquini et al. (2023), have advanced our understanding of its structure. However, the gluon and sea-quark content remains less constrained than in the nucleon due to limited experimental data. Upcoming facilities like Jefferson Lab Arrington et al. (2022) and future electron-ion colliders (EICs) Abdul Khalek et al. (2022); Anderle et al. (2021) aim to probe pion PDFs via the Sullivan process Sullivan (1972). The COMPASS++/AMBER program will further explore pion structure using high-intensity pion beams Adams et al. (2018).
@@
In this work, we develop a light-front model of the pion that explicitly incorporates gluonic degrees of freedom, providing a framework to explore the pion’s gluonic structure. In this approach, high-energy scattering processes involving the pion are modeled as being mediated by an active gluon, while the remaining constituents form an effective spectator system characterized by a dynamically generated mass. The mass and LFWFs of the pion are determined self-consistently by solving two complementary equations rooted in QCD: the light-front holographic Schrödinger equation, which captures the transverse dynamics and confinement mechanism in the chiral limit Brodsky et al. (2015), and the ’t Hooft equation, which governs the longitudinal momentum structure of mesons in the large-$N_{c}$ limit ’t Hooft (1974). Together, these equations offer a three-dimensional description of the pion’s bound-state dynamics, incorporating both the transverse and longitudinal aspects of confinement. This unified framework provides predictions for gluon-sensitive observables, including the gluon PDFs, GPDs, TMDs, and GFFs of the pion. The model predictions show good agreement with recent lattice QCD results and the available global fits, demonstrating its effectiveness in capturing both the momentum and spatial structure of the pion.
@@ II Theoretical framework
Table 1: Quantum numbers and masses of the pion family.
@@ III Pion Mass Spectroscopy and Light-Front Wavefunctions
Although the longitudinal wavefunction $\chi(x)$ cannot be determined analytically, it can be well approximated by:
@@
$|\pi(P)\rangle=$ $\int\frac{\mathrm{d}^{2}\vec{k}_{\perp}\mathrm{d}x}{16\pi^{3}\sqrt{x(1-x)}}\Big{[}\psi_{-1,+1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|-1,+1;xP^{+},\vec{k}_{\perp}\right\rangle$
$+\psi_{+1,-1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|+1,-1;xP^{+},\vec{k}_{\perp}\right\rangle\Big{]}.$ (15)
@@ IV Gluon generalized parton distributions
$H^{g}_{\pi}(x,\,\xi,\,t)=\frac{1}{\bar{P}^{+}}\int\frac{dz^{-}}{2\pi}e^{ix\bar{P}^{+}z^{-}}\langle\pi(P^{\prime})|\,F^{+i}\left(-\frac{z}{2}\right)F^{+i}\left(\frac{z}{2}\right)\,|\pi(P)\rangle\Big{|}_{\begin{subarray}{c}z^{+}=0\\ z^{\perp}=0\end{subarray}},$ (17)
@@
$\left[{\rm d}\mathcal{X}\,{\rm d}\mathcal{K}_{\perp}\right]=$ $\sqrt{x^{2}-\xi^{2}}(\sqrt{1-\xi^{2}}\ )^{-1}\prod_{i=1}^{2}\left[\frac{{\rm d}x_{i}{\rm d}^{2}\vec{k}_{i\perp}}{16\pi^{3}}\right]$
×$\times\delta(x-x_{1})16\pi^{3}\delta\left(1-\sum_{i=1}^{2}x_{i}\right)\delta^{2}\left(\sum_{i=1}^{2}\vec{k}_{i\perp}\right),$ (21)
and the light-front momenta are $x^{\prime}_{1}=\frac{x_{1}+\xi}{1+\xi}$; $\vec{k}^{\prime}_{1\perp}=\vec{k}_{1\perp}+(1-x^{\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the initial gluon ($i=1$) and $x^{\prime}_{i}=\frac{x_{i}}{1+\xi};~{}\vec{k}^{\prime}_{i\perp}=\vec{k}_{i\perp}-{x_{i}^{\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the initial spectator ($i\neq 1$), and $x^{\prime\prime}_{1}=\frac{x_{1}-\xi}{1-\xi}$; $\vec{k}^{\prime\prime}_{1\perp}=\vec{k}_{1\perp}-(1-x^{\prime\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the final gluon and $x^{\prime\prime}_{i}=\frac{x_{i}}{1-\xi};~{}\vec{k}^{\prime\prime}_{i\perp}=\vec{k}_{i\perp}+{x_{i}^{\prime\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the final spectator. Note that, in this work, we limit ourselves to the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) region, $\xi<x<1$, where the number of partons in the initial and the final states remains conserved.
Figure 3: The gluon GPD of the pion as a function of (a) $x$ and $-t$ [in GeV2] at fixed $\xi$, and (b) $x$ and $\xi$ at fixed $-t$ [in GeV2].
@@
Figure 4: The gluon PDF of the pion (a) at the model scale, and (b) at the scale $5$ GeV2 compared with the available global fits Barry et al. (2021); Novikov et al. (2020); Pasquini et al. (2023).
@@ IV.1 Gluon PDF
Figure 5: The gluon GPD of the pion as a function of (a) $x$ and ${b_{\perp}}$ [in fm] at fixed $\xi$, and (b) $\xi$ and ${b_{\perp}}$ [in fm] at fixed $x$.
@@ IV.2 Gluon distributions in transverse position space
$H^{g}_{\pi}(x,\xi,{b}_{\perp})\;=\;\frac{1}{\left(1-\xi^{2}\right)}\int\frac{d^{2}\vec{\Delta}_{\perp}}{(2\pi)^{2}}\,e^{-i\frac{\vec{\Delta}_{\perp}}{\left(1-\xi^{2}\right)}\cdot\vec{b}_{\perp}}\,H^{g}_{\pi}\!\left(x,\xi,t=-\frac{\Delta_{\perp}^{2}+4\xi^{2}M^{2}}{1-\xi^{2}}\right),$ (22)
@@ IV.3 Gluon distributions in boost-invariant longitudinal position space
$H(x,\sigma,t)=\int_{0}^{\xi_{\rm f}}\frac{d\xi}{2\pi}\,e^{i\xi\sigma}H(x,\xi,t),$ (23)
@@
We also present the gluon GPDs in longitudinal position space, $H(x,\sigma,t)$, as a function of the longitudinal impact parameter $\sigma$, for a fixed value of momentum transfer squared $t=-0.5~{}\text{GeV}^{2}$ and at different values of the longitudinal momentum fraction $x$ in the right panel of Fig. 6. As $x$ increases, the diffraction pattern becomes more pronounced, with sharper peaks and a more localized distribution in $\sigma$. This indicates that gluons carrying a larger fraction of the pion’s longitudinal momentum are more tightly localized in longitudinal position space. Conversely, for smaller $x$, the distribution appears broader, reflecting a more delocalized gluonic structure. This behavior highlights the strong correlation between the momentum fraction and the spatial location of the partons encoded in the GPDs. Note that similar features have also been observed in the gluon GPDs of the proton Zhang et al. (2025), as well as in DVCS amplitudes, coordinate-space parton distributions, and Wigner distributions across a range of models Brodsky et al. (2006, 2007); Miller and Brodsky (2020); Chakrabarti et al. (2009); Manohar et al. (2011); Kumar and Dahiya (2015); Mondal and Chakrabarti (2015); Chakrabarti and Mondal (2015); Mondal (2017); Kaur et al. (2018); Maji et al. (2022); Ojha et al. (2023).
Figure 7: The GFF of pion with respect to $Q^{2}$ [in GeV2].
@@ IV.4 Pion gravitational form factor
where $Q^{2}=\Delta_{\perp}^{2}$. Figure 7 shows the GFF $A(Q^{2})$ of the pion, including separate contributions from the gluon and the spectator. At the model scale, we observe that the spectator contribution strongly dominates over the gluonic contribution. We compare our total result, $A_{\pi}(Q^{2})=A^{g}_{\pi}(Q^{2})+A^{s}_{\pi}(Q^{2})$, with the latest lattice QCD simulations Hackett et al. (2023) and find good agreement. From this result, we extract the pion’s mass (or matter) radius as $r^{2}_{\rm m}=-6A^{\prime}(0)/A(0)=(0.46~{}\mathrm{fm})^{2}$, which is very close to the corresponding lattice QCD value $r^{2}_{\rm m}=(0.41~{}\mathrm{fm})^{2}$ Hackett et al. (2023).
@@ V Transverse momentum-dependent gluon distributions
The gluon TMDs can be expressed in terms of overlaps of LFWFs as follows
$f_{1\pi}^{g}(x,{k}^{2}_{\perp})=\sum_{\lambda_{g},\lambda_{s}}\int\frac{{\rm d}x~{}{\rm d}\vec{k}_{\perp}}{16\pi^{3}}\Big{|}\psi^{\pi}_{\lambda_{g},\lambda_{s}}(x,\vec{k}_{\perp})\Big{|}^{2}.$ (27)
Figure 8: The $x$ multiplied TMD of gluon in pion with respect to $x$ and $k^{2}_{\perp}$ [in GeV2].
@@
Figure 9: The first and second $k_{\perp}$-moment of gluon TMD in pion with respect to $x$.
@@
$\langle k_{\perp}\rangle^{g}_{\pi}=0.175~{}\text{GeV},\qquad\langle k_{\perp}^{2}\rangle^{g}_{\pi}=0.043~{}\text{GeV}^{2},$

### v2

@@ I Introduction
Global analyses of pion PDFs, primarily using Drell-Yan, $J/\psi$ production, and direct photon data Barry et al. (2021); Novikov et al. (2020); Pasquini et al. (2023); Kotz et al. (2025), have advanced our understanding of its structure. However, the gluon and sea-quark content remains less constrained than in the nucleon due to limited experimental data. Upcoming facilities like Jefferson Lab Arrington et al. (2022) and future electron-ion colliders (EICs) Abdul Khalek et al. (2022); Anderle et al. (2021) aim to probe pion PDFs via the Sullivan process Sullivan (1972). The COMPASS++/AMBER program will further explore pion structure using high-intensity pion beams Adams et al. (2018).
@@
In this work, we develop a light-front model of the pion that explicitly incorporates gluonic degrees of freedom, providing a framework to explore the pion’s gluonic structure. In this approach, high-energy scattering processes involving the pion are modeled as being mediated by an active gluon, while the remaining constituents form an effective spectator system with an effective mass, motivated by the gluon-spectator model for the nucleon Lyubovitskij and Schmidt (2021); Chakrabarti et al. (2023). The mass and LFWFs of the pion are determined self-consistently by solving two complementary equations rooted in QCD: the light-front holographic Schrödinger equation, which captures the transverse dynamics and confinement mechanism in the chiral limit Brodsky et al. (2015), and the ’t Hooft equation, which governs the longitudinal momentum structure of mesons in the large-$N_{c}$ limit ’t Hooft (1974). Together, these equations offer a three-dimensional description of the pion’s bound-state dynamics, incorporating both the transverse and longitudinal aspects of confinement. This unified framework provides predictions for gluon-sensitive observables, including the gluon PDFs, GPDs, TMDs, and GFFs of the pion. The model predictions show good agreement with recent lattice QCD results and the available global fits, demonstrating its effectiveness in capturing both the momentum and spatial structure of the pion.
@@ II Theoretical framework
Table 1: Quantum numbers and masses of the pion family. The transverse mass $M_{\perp}$ and longitudinal mass $M_{\parallel}$ are obtained by solving the holographic Schrödinger equation and the ’t Hooft equation, respectively. The total mass $M$ of the system is then computed using Eq. (11).
@@ III Pion Mass Spectroscopy and Light-Front Wavefunctions
Although the longitudinal mode $\chi(x)$ cannot be determined analytically, it can be well approximated by:
@@
$|\pi(P)\rangle=$ $\int\frac{\mathrm{d}x\,\mathrm{d}^{2}\vec{k}_{\perp}}{16\pi^{3}\sqrt{x(1-x)}}\Big{[}\psi_{-1,+1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|-1,+1;xP^{+},\vec{k}_{\perp}\right\rangle+\psi_{+1,-1}^{\pi}\left(x,\vec{k}_{\perp}\right)\left|+1,-1;xP^{+},\vec{k}_{\perp}\right\rangle\Big{]}.$ (15)
@@ IV Gluon generalized parton distributions
$H^{g}_{\pi}(x,\,\xi,\,t)=\frac{1}{\bar{P}^{+}}\int\frac{{\rm d}z^{-}}{2\pi}e^{ix\bar{P}^{+}z^{-}}\langle\pi(P^{\prime})|\,F^{+i}\left(-\frac{z}{2}\right)F^{+i}\left(\frac{z}{2}\right)\,|\pi(P)\rangle\Big{|}_{\begin{subarray}{c}z^{+}=0\\ z^{\perp}=0\end{subarray}},$ (17)
@@
$\left[{\rm d}\mathcal{X}\,{\rm d}\mathcal{K}_{\perp}\right]\equiv$ $\sqrt{x^{2}-\xi^{2}}(\sqrt{1-\xi^{2}}\ )^{-1}\prod_{i=1}^{2}\left[\frac{{\rm d}x_{i}{\rm d}^{2}\vec{k}_{i\perp}}{16\pi^{3}}\right]\delta(x-x_{1})16\pi^{3}\delta\left(1-\sum_{i=1}^{2}x_{i}\right)\delta^{2}\left(\sum_{i=1}^{2}\vec{k}_{i\perp}\right),$
and the light-front momenta are $x^{\prime}_{1}=\frac{x_{1}+\xi}{1+\xi}$; $\vec{k}^{\prime}_{1\perp}=\vec{k}_{1\perp}+(1-x^{\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the initial gluon ($i=1$) and $x^{\prime}_{i}=\frac{x_{i}}{1+\xi};\leavevmode\nobreak\ \vec{k}^{\prime}_{i\perp}=\vec{k}_{i\perp}-{x_{i}^{\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the initial spectator ($i\neq 1$), and $x^{\prime\prime}_{1}=\frac{x_{1}-\xi}{1-\xi}$; $\vec{k}^{\prime\prime}_{1\perp}=\vec{k}_{1\perp}-(1-x^{\prime\prime})\frac{\vec{\Delta}_{\perp}}{2}$ for the final gluon and $x^{\prime\prime}_{i}=\frac{x_{i}}{1-\xi};\leavevmode\nobreak\ \vec{k}^{\prime\prime}_{i\perp}=\vec{k}_{i\perp}+{x_{i}^{\prime\prime}}\frac{\vec{\Delta}_{\perp}}{2}$ for the final spectator. Note that, in this work, we limit ourselves to the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) region, $\xi<x<1$, where the number of partons in the initial and the final states remains conserved.
Figure 3: The gluon GPD of the pion as a function of (a) $x$ and $-t$ [in GeV2] at fixed skewness, $\xi=0.15$, and (b) $x$ and $\xi$ at fixed $-t=0.5$ GeV2.
@@ IV.1 Gluon PDF
[…]
The right panel of Fig. 4 shows our results for the pion’s gluon PDF after QCD evolution. The error bands reflect a 10% uncertainty in the initial scale. As shown, our results show similar behavior to existing global fits Barry et al. (2021); Novikov et al. (2020); Kotz et al. (2025) and phenomenological predictions Pasquini et al. (2023); Lan et al. (2025), with better agreement observed with the JAM21 global QCD analysis Barry et al. (2021).
@@ IV.2 Gluon distributions in transverse position space
$H^{g}_{\pi}(x,\xi,{b}_{\perp})\;=\;\frac{1}{\left(1-\xi^{2}\right)}\int\frac{{\rm d}^{2}\vec{\Delta}_{\perp}}{(2\pi)^{2}}\,e^{-i\frac{\vec{\Delta}_{\perp}}{\left(1-\xi^{2}\right)}\cdot\vec{b}_{\perp}}\,H^{g}_{\pi}\!\left(x,\xi,t=-\frac{\Delta_{\perp}^{2}+4\xi^{2}M^{2}}{1-\xi^{2}}\right),$ (21)
@@
Figure 5: The gluon GPD of the pion in transverse position space as a function of (a) $x$ and $b_{\perp}$ at fixed skewness, $\xi=0.15$, and (b) $\xi$ and $b_{\perp}$ at fixed $x=0.6$.
@@ IV.3 Gluon distributions in boost-invariant longitudinal position space
$H(x,\sigma,t)=\int_{0}^{\xi_{\rm f}}\frac{{\rm d}\xi}{2\pi}\,e^{i\xi\sigma}H(x,\xi,t),$ (22)
@@
We also present the gluon GPDs in longitudinal position space, $H(x,\sigma,t)$, as a function of the longitudinal impact parameter $\sigma$, for a fixed value of momentum transfer squared $t=-0.5\leavevmode\nobreak\ \text{GeV}^{2}$ and at different values of the longitudinal momentum fraction $x$ in the right panel of Fig. 6. As $x$ increases, the diffraction pattern becomes more pronounced, with sharper peaks and a more localized distribution in $\sigma$. This indicates that gluons carrying a larger fraction of the pion’s longitudinal momentum are more tightly localized in longitudinal position space. Conversely, for smaller $x$, the distribution appears broader, reflecting a more delocalized gluonic structure. This behavior highlights the strong correlation between the momentum fraction and the spatial location of the partons encoded in the GPDs. Note that similar features have also been observed in the gluon GPDs of the proton Zhang et al. (2025), as well as in DVCS amplitudes, coordinate-space parton distributions, and Wigner distributions across a range of models Brodsky et al. (2006, 2007); Miller and Brodsky (2020); Chakrabarti et al. (2009); Manohar et al. (2011); Kumar and Dahiya (2015); Mondal and Chakrabarti (2015); Chakrabarti and Mondal (2015); Mondal (2017); Kaur et al. (2018); Maji et al. (2022); Ojha et al. (2023).
Figure 7: Our prediction for the GFF of the pion, showing separate contributions from the gluon (black dashed line) and the spectator (orange dot-dashed line): (a) $A(Q^{2})$, and (b) $Q^{2}A(Q^{2})$. The total GFF (blue solid line) is compared with the lattice QCD results Hackett et al. (2023).
@@ IV.4 Pion gravitational form factor
where $Q^{2}=\Delta_{\perp}^{2}$. Figure 7 shows the GFF $A(Q^{2})$ of the pion, including separate contributions from the gluon and the spectator. At the model scale, we observe that the spectator contribution strongly dominates over the gluonic contribution. We compare our total result, $A_{\pi}(Q^{2})=A^{g}_{\pi}(Q^{2})+A^{s}_{\pi}(Q^{2})$, with the latest lattice QCD simulations Hackett et al. (2023) and find good agreement. From this result, we extract the pion’s mass (or matter) radius as $r^{2}_{\rm m}=-6A^{\prime}(0)/A(0)=(0.46\leavevmode\nobreak\ \mathrm{fm})^{2}$, which is very close to the corresponding lattice QCD value $r^{2}_{\rm m}=(0.41\leavevmode\nobreak\ \mathrm{fm})^{2}$ Hackett et al. (2023).
@@ V Transverse momentum-dependent gluon distributions
The gluon TMDs can be expressed in terms of overlaps of LFWFs as follows Chakrabarti et al. (2023)
$f_{1\pi}^{g}(x,{k}^{2}_{\perp})=\frac{1}{16\pi^{3}}\sum_{\lambda_{g},\lambda_{s}}\left|\psi^{\pi}_{\lambda_{g},\lambda_{s}}(x,\vec{k}_{\perp})\right|^{2}.$ (26)
Figure 8: The $x$-weighted gluon TMD, $f_{1\pi}^{g}(x,k_{\perp}^{2})$, in the pion as a function of $x$ and $k_{\perp}^{2}$.
@@
Figure 9: The first and second $k_{\perp}$-moments of the gluon TMD $f_{1\pi}^{g}(x,k_{\perp}^{2})$ in the pion, shown as functions of $x$.
@@
$\langle k_{\perp}\rangle^{g}_{\pi}=0.175\leavevmode\nobreak\ \text{GeV},\qquad\langle k_{\perp}^{2}\rangle^{g}_{\pi}=0.043\leavevmode\nobreak\ \text{GeV}^{2},$